OpenStudy (anonymous):

is it possible to find the nth term for the series 1,5,7,11,13,17,19,....

OpenStudy (anonymous):

no, i came across this series when i had to find the nth harmonics in motor current excluding the even harmonics and multiples of 3 harmonics

OpenStudy (anonymous):

Well its pretty vague, you can't determine the nth term of a sequence of terms just by giving a few of them

OpenStudy (anonymous):

You can guess tho lol

OpenStudy (anonymous):

well , 6n+1 gives the series 1,7,13,.... and 6n-1 gives 5,11,17,.... i want them both in a single expression

OpenStudy (anonymous):

Then you should have just asked for that in the first place.

OpenStudy (anonymous):

well i m sorry... i was trying to explain that i didn't them from working on prime numbers

OpenStudy (anonymous):

You want $$\{ 6n-1 , n\in \mathbb{N}\}\bigcup \{ 6n+1 , n\in \mathbb{N}\}$$

OpenStudy (anonymous):

you got it.... but how do i put it as a single expression?

OpenStudy (anonymous):

I just did

OpenStudy (anonymous):

a single expression which can be put as a statement in a program.... which will give the series 1,5,7,11,13,.... for each successive value of n (where n is a whole number)

OpenStudy (anonymous):

k one sec

OpenStudy (zzr0ck3r):

it seems your asking for a solution to 6n+1=6n-1

OpenStudy (anonymous):

take your time buddy

OpenStudy (anonymous):

@abhinayreddy here it is i got it

OpenStudy (anonymous):

my bad one sec

OpenStudy (zzr0ck3r):

that worked, you just had to start at n = -1, very cool....

OpenStudy (anonymous):

@abhinayreddy $$\frac{1}{2}(6n-3+(-1)^n)$$

OpenStudy (anonymous):

@jack119how did you arrive at that expression.....?

OpenStudy (anonymous):

please explain me.... and thanks a lot man